# Deconstructing The Ethics of Large Language Models from Long-standing Issues to New-emerging Dilemmas

### v1

@@ 2.3.1 Mitigation Strategy
Pre-LLM Era.
As machine learning models are increasingly deployed in critical domains [144, 306, 346, 117, 345], addressing bias to achieve fairness has become essential. Traditional bias mitigation approaches are categorized into three main strategies. Pre-processing techniques aim to modify the data by reducing inherent biases [62]. For example, Pessach et al. [225] suggest a pre-processing mechanism to enhance fairness in private collaborative machine learning scenarios [334, 48]. In-processing methods involve altering learning algorithms to eliminate bias during model training [303]. Berk et al. [25] introduced fairness regularizers for linear and logistic regression models to ensure both group and individual fairness. Post-processing techniques are applied after training, adjusting model outputs to enhance fairness [142]. Petersen et al. [226] developed a general post-processing algorithm that ensures individual fairness by utilizing graph Laplacian regularization [292], framing the challenge as a graph smoothing problem.
LLM Era.
Bias mitigation techniques in LLMs also follow a similar pattern and can be categorized into four groups based on their application at different stages of the LLM workflow: pre-processing, in-training, intra-processing, and post-processing [77].
@@ 2.3.2 Measurements on Fairness
Discussion.
To effectively mitigate bias in LLMs, it is essential to adopt a comprehensive approach that leverages the strengths of various bias mitigation strategies. Specifically, pre-processing techniques should be employed to neutralize biases at the source, ensuring that the data used to train the LLM is as unbiased as possible. Subsequently, in-training mitigation strategies can be implemented to further refine the training process of the LLM, improving its ability to produce fair and unbiased outputs. Finally, during the model’s deployment phase, both intra-processing and post-processing measures could be applied to minimize the risk of generating biased content. By combining these methods, we can create a robust framework that significantly reduces the likelihood of bias in outputs, fostering a more equitable and fair use of LLMs.
@@ 3.1.1 Hallucination
Large language models tend to produce hallucinations where the models generate contents that deviate from the input, contradict existing contexts, or misalign with universally accepted world knowledge [153, 282, 314, 83]. Such phenomena pose significant challenges, particularly when considering the reliability and trustworthiness of LLMs in critical applications. We delve into the underlying causes, manifestations, and potential mitigation strategies for hallucinations in LLMs.
Underlying Causes.
The primary causes of hallucinations in LLMs can be broadly categorized into data quality, model architecture, and algorithmic limitations:
@@
Manifestations.
Hallucinations in LLMs manifest in various forms, from minor inaccuracies to entirely fictitious narratives. Sometimes, these manifest as confident but false assertions, particularly misleading when LLMs are employed in sensitive fields such as medical diagnostics [131], legal advising [213], social content moderation [211], or education [243].
Mitigation Strategies.
Numerous research has attempted to mitigate hallucination in LLMs [119]. Most existing mitigation strategies can be categorized into data-centric approaches [204, 89, 3, 234, 139, 341] and model-centric approaches [150, 160, 218, 276]. In the data-centric approaches, several works aim to improve the quality of training data, ensuring it is accurate, diverse, and free of biases. This may involve rigorous data curation and validation processes [167]. Tian et al. introduce the external knowledge graph to mitigate the problem of hallucinations [262]. For the model-centric approaches, many works enhance the model architectures for a better understanding of context, discern factual accuracy, and recognize when the model is venturing into areas of low confidence or outside its training scope [133]. This could involve incorporating mechanisms to check factual accuracy in real time or integrating feedback loops that allow the model to learn from its mistakes. Yao et al. directly edit model parameters to bridge the knowledge gap to mitigate hallucinations [311]. While substantial progress has been made in identifying and categorizing hallucinations [153], the development of robust mechanisms to prevent or correct these errors remains an ongoing area of research. This is crucial for LLMs’ future advancements in various fields.
Discussion.
Detecting instances when LLMs are prone to hallucinations is crucial. While the bulk of research on LLM hallucination has centered on the English language, it has been shown that these models are more prone to hallucinations in non-English languages [131]. This disparity underscores a significant gap in our understanding of hallucinations within multilingual contexts and underscores the urgency in developing robust detection and mitigation strategies for hallucinations in diverse linguistic environments. Furthermore, most existing studies have been centered around unimodal hallucinations. However, the emergence of multimodal LLMs, capable of synthesizing and interpreting data across different modalities such as text, images, and audio, poses unique challenges [274, 167, 79, 81, 312]. Overall, addressing hallucination effectively in LLMs requires a comprehensive approach that encompasses multiple languages, modalities, and cultural contexts. Furthermore, transparency regarding operational mechanisms and the inherent limitations of models is vital. Educating users about the potential for hallucinations and the specific contexts in which they are most likely to occur can enable a more critical evaluation of outputs generated by LLMs.
@@ 3.1.2 Sycophancy
Large language models may exhibit a tendency to flatter users by reaffirming their misconceptions and stated beliefs, a behavior known as sycophancy [121]. This issue raises significant concerns about the models’ ability to provide objective and unbiased information. Sycophancy in LLMs can lead to the reinforcement of incorrect beliefs, limiting the educational and corrective potential of these systems, and potentially exacerbating echo chambers in digital interactions [247, 147].
Underlying Causes.
The propensity for sycophancy can be attributed to several factors:
@@
Discussion.
Future research directions to investigate and resolve the issue of sycophancy in LLMs should focus on several key areas. Firstly, developing methods for detecting when an LLM is likely to be reinforcing misconceptions is crucial. This involves enhancing the model’s ability to recognize and differentiate between fact-based assertions and user opinions. Secondly, there is a need to design algorithms that can introduce a balance between user engagement and factual integrity. These algorithms would ensure that while user interactions remain engaging, they do not compromise on delivering accurate and unbiased information. Moreover, exploring the implementation of feedback mechanisms where users can flag responses perceived as overly agreeable or flattering could provide valuable data for training more objective models. Lastly, interdisciplinary research incorporating insights from psychology and ethics could guide the development of LLMs that maintain a neutral stance, particularly in sensitive or polarized topics. These efforts are essential for advancing LLM technology to be both useful and ethically responsible.
@@ 3.2 Social Norm
What is alignment in LLMs and why is it needed?
With the transformative evolution in Natural Language Processing (NLP) research and development, the impact and success of large language models (LLMs) [332, 330, 57, 323, 261, 5, 268, 269] has been exceptional, exemplified by ChatGPT [298] developed by OpenAI. One key driver for the popularity and usability of recent LLMs is alignment. Alignment is a technique that aims to ensure that generated responses comply with human values. Currently, the standard procedure for aligning large language models (LLMs) primarily includes two approaches: SFT (Supervised Fine-Tuning) [217] and RLHF (Reinforcement Learning from Human Feedback) [55, 17]. Since LLMs have been used in a wide range of applications (e.g., editing/writing assistance, personal consultation, question answering, and customer support), many corresponding concerns would arise if the LLMs are not properly aligned otherwise.
@@
To understand the characteristics of LLMs, a diverse array of benchmarks have been introduced [163, 279, 277, 278]. In contrast to general-purpose evaluation, alignment-focused evaluation depends on the taxonomy of alignment, associated with corresponding scenarios, criteria, and datasets [281, 344, 343]. Obtaining appropriate criteria and datasets for evaluating alignments in LLMs is crucial, albeit a non-trivial task  [51, 278]. This essentially involves representing the preferences of humans [41]. However, manually collecting human judgment can be expensive, time-consuming, and labor-intensive [339]. To address this issue, researchers proposed to use strong LLMs as an automated proxy for evaluating other LLMs [342]. For example, AUTO-J [152] is trained to tackle challenges in evaluating LLM alignments regarding generality, flexibility, and interoperability. AUTOCALIBRATE presents a multi-stage, gradient-free approach [152], to automatically calibrate and align an LLM-based evaluator toward human preference free of human intervention.
Recent Advancements in Alignment.
@@
In the endeavor to align LLMs with human values, a myriad of research initiatives [258, 240, 252, 297, 307, 128, 244] have been undertaken to achieve effective LLM alignments. The forefront of these approaches emphasizes the generative capabilities of large language models (LLMs) for self-regulation with minimal human supervision. SELF-ALIGN [258] proposes a topic-guided, principle-driven approach to autonomously generate responses that are helpful, ethical, and reliable, leveraging the mechanism of in-context learning. Similarly, KNOWNO [240] is a framework for evaluating and aligning the uncertainty in LLM-based planning. Utilizing the theory of conformal prediction, KNOWNO ensures statistical reliability in task completion, thereby minimizing human assistance in complex planning scenarios. Additionally, PRO [252] introduces a response probability ranking method, enhancing the Bradley-Terry comparison model to effectively direct the LLM to favor the most appropriate response. Complementarily, P3O [297] presents a trajectory-wise policy gradient algorithm, which uniquely focuses on comparative rewards instead of traditional reward optimization trained from comparison-based losses.
Discussion.
@@
The burgeoning field of LLM alignment, pivotal for the symbiosis of AI and humanity, anticipates transformative discoveries. Emphasizing the importance of AI safety and the seamless integration of AI with human society, prioritizing the alignment of LLMs, with human ethos is essential. As LLMs’ capabilities escalate, the complexity of achieving this alignment intensifies, necessitating increased scientific and technological investment. This demands an exploration of novel strategies in this domain. Foremost, amidst the rapid evolution of LLMs, it is crucial to guarantee their adherence to human ethical standards, which requires more theoretical breakthroughs [296]. In addition, the growing intricacy of AI architectures calls for automated systems capable of assessing and realigning these models [223]. Next, the black-box nature of LLMs also highlights the urgency for clarity and explainability in their alignment processes [337]. Lastly, leveraging adversarial attacks as a method to test and refine the alignment of LLMs emerges as an effective approach for ensuring their conformity to human values [349].
@@ 3.3 Law and Regulatory Compliance
Discussion.
Despite the heroic striving of the AI Act to keep up with the accelerating dynamics of AI development, several discussions are also proposed around its practical compliance with LLMs. Hacker et al. argue that this direct regulation is unsatisfactory and could be further enhanced from 1) the definition of GPAIS, 2) the risk management of GPAIS, and 3) the adverse consequences for competition [94]. They propose to focus on the deployers and users more and directly apply non-discrimination and data protection law (GDPR compliance) on LLMs. Bommasani et al. [34] systematically evaluate the compliance with the draft EU AI Act of the foundation model providers like OpenAI and Google. They evaluate the compliance of 10 major foundation model providers (and their flagship models) with the 12 requirements proposed by the EU AI Act and use a scale from 0 (worst) to 4 (best) to rate each provider and model for each requirement. The best possible score for a provider or a model is 48, which indicates full compliance with the AI Act. Their results identify four areas where many organizations receive low scores (usually 0 or 1 out of 4) in terms of compliance with the AI Act: 1) copyrighted data, 2 compute/energy, 3) risk mitigation, and 4) evaluation/testing. Aside from these general regulations, there are also discussions on challenges of how to regulate LLMs for vertical domains such as medical usage [200] and healthcare [199].

### v2

@@ 2.3.1 Mitigation Strategy
Bias mitigation in traditional machine learning involves pre-processing data to reduce bias, altering algorithms during training (in-processing), and adjusting outputs post-training (post-processing). In the LLM era, similar strategies are employed: pre-processing techniques reduce bias in training data and prompts, in-training methods modify training procedures and model architecture, intra-processing approaches generate debiased predictions during inference, and post-processing techniques address bias in outputs, particularly for black-box models.
Pre-LLM Era. As machine learning models are increasingly deployed in critical domains [145, 311, 352, 118, 351], addressing bias to achieve fairness has become essential. Traditional bias mitigation approaches are categorized into three main strategies. Pre-processing techniques aim to modify the data by reducing inherent biases [63]. For example, Pessach et al. [229] suggest a pre-processing mechanism to enhance fairness in private collaborative machine learning scenarios [340, 48]. In-processing methods involve altering learning algorithms to eliminate bias during model training [308]. Berk et al. [25] introduced fairness regularizers for linear and logistic regression models to ensure both group and individual fairness. Post-processing techniques are applied after training, adjusting model outputs to enhance fairness [143]. Petersen et al. [230] developed a general post-processing algorithm that ensures individual fairness by utilizing graph Laplacian regularization [297], framing the challenge as a graph smoothing problem.
LLM Era. Bias mitigation techniques in LLMs also follow a similar pattern and can be categorized into four groups based on their application at different stages of the LLM workflow: pre-processing, in-training, intra-processing, and post-processing [78].
@@ 2.3.2 Measurements on Fairness
Discussion. To effectively mitigate bias in LLMs, it is essential to adopt a comprehensive approach that leverages the strengths of various bias mitigation strategies. Specifically, pre-processing techniques should be employed to neutralize biases at the source, ensuring that the data used to train the LLM is as unbiased as possible. Subsequently, in-training mitigation strategies can be implemented to further refine the training process of the LLM, improving its ability to produce fair and unbiased outputs. Finally, during the model’s deployment phase, both intra-processing and post-processing measures could be applied to minimize the risk of generating biased content. By combining these methods, we can create a robust framework that significantly reduces the likelihood of bias in outputs, fostering a more equitable and fair use of LLMs.
@@ 3.1.1 Hallucination
Large language models tend to produce hallucinations where the models generate contents that deviate from the input, contradict existing contexts, or misalign with universally accepted world knowledge [155, 287, 320, 84]. An example is shown in Figure 6. Such phenomena pose significant challenges, particularly when considering the reliability and trustworthiness of LLMs in critical applications. We delve into the underlying causes, manifestations, and potential mitigation strategies for hallucinations in LLMs.
Figure 6: An example of hallucination. The LLM correctly identifies Thomas Edison as a key figure in the invention of the lightbulb in 1879. However, the model also fabricates information about Benjamin Franklin inventing an electric light in 1802, which is inaccurate. Benjamin Franklin is well-known for his experiments with electricity, particularly the kite experiment, but he did not contribute to the invention of the lightbulb. The model ”hallucinated” this fact, likely by confusing Franklin’s work with electricity with the development of the lightbulb.
Underlying Causes. The primary causes of hallucinations in LLMs can be broadly categorized into data quality, model architecture, and algorithmic limitations:
@@
Manifestations. Hallucinations in LLMs manifest in various forms, from minor inaccuracies to entirely fictitious narratives. Sometimes, these manifest as confident but false assertions, particularly misleading when LLMs are employed in sensitive fields such as medical diagnostics [132], legal advising [217], social content moderation [215], or education [247].
Mitigation Strategies. Numerous research has attempted to mitigate hallucination in LLMs [120]. Most existing mitigation strategies can be categorized into data-centric approaches [208, 90, 3, 238, 140, 347] and model-centric approaches [151, 162, 222, 281]. In the data-centric approaches, several works aim to improve the quality of training data, ensuring it is accurate, diverse, and free of biases. This may involve rigorous data curation and validation processes [169]. Tian et al. introduce the external knowledge graph to mitigate the problem of hallucinations [267]. For the model-centric approaches, many works enhance the model architectures for a better understanding of context, discern factual accuracy, and recognize when the model is venturing into areas of low confidence or outside its training scope [134]. This could involve incorporating mechanisms to check factual accuracy in real time or integrating feedback loops that allow the model to learn from its mistakes. Yao et al. directly edit model parameters to bridge the knowledge gap to mitigate hallucinations [317]. While substantial progress has been made in identifying and categorizing hallucinations [155], the development of robust mechanisms to prevent or correct these errors remains an ongoing area of research. This is crucial for LLMs’ future advancements in various fields.
Discussion. Detecting instances when LLMs are prone to hallucinations is crucial. While the bulk of research on LLM hallucination has centered on the English language, it has been shown that these models are more prone to hallucinations in non-English languages [132]. This disparity underscores a significant gap in our understanding of hallucinations within multilingual contexts and underscores the urgency in developing robust detection and mitigation strategies for hallucinations in diverse linguistic environments. Furthermore, most existing studies have been centered around unimodal hallucinations. However, the emergence of multimodal LLMs, capable of synthesizing and interpreting data across different modalities such as text, images, and audio, poses unique challenges [279, 169, 80, 82, 318, 153]. Overall, addressing hallucination effectively in LLMs requires a comprehensive approach that encompasses multiple languages, modalities, and cultural contexts. Furthermore, transparency regarding operational mechanisms and the inherent limitations of models is vital. Educating users about the potential for hallucinations and the specific contexts in which they are most likely to occur can enable a more critical evaluation of outputs generated by LLMs.
@@ 3.1.2 Sycophancy
Large language models may exhibit a tendency to flatter users by reaffirming their misconceptions and stated beliefs, a behavior known as sycophancy [122]. This issue raises significant concerns about the model’s ability to provide objective and unbiased information. Sycophancy in LLMs can lead to the reinforcement of incorrect beliefs, limiting the educational and corrective potential of these systems, and potentially exacerbating echo chambers in digital interactions [251, 148].
Underlying Causes. The propensity for sycophancy can be attributed to several factors:
@@
Discussion. Future research directions to investigate and resolve the issue of sycophancy in LLMs should focus on several key areas. Firstly, developing methods for detecting when an LLM is likely to be reinforcing misconceptions is crucial. This involves enhancing the model’s ability to recognize and differentiate between fact-based assertions and user opinions. Secondly, there is a need to design algorithms that can introduce a balance between user engagement and factual integrity. These algorithms would ensure that while user interactions remain engaging, they do not compromise on delivering accurate and unbiased information. Moreover, exploring the implementation of feedback mechanisms where users can flag responses perceived as overly agreeable or flattering could provide valuable data for training more objective models. Lastly, interdisciplinary research incorporating insights from psychology and ethics could guide the development of LLMs that maintain a neutral stance, particularly in sensitive or polarized topics. These efforts are essential for advancing LLM technology to be both useful and ethically responsible.
@@ 3.2 Social Norm
What is alignment in LLMs and why is it needed? With the transformative evolution in Natural Language Processing (NLP) research and development, the impact and success of large language models (LLMs) [338, 336, 57, 329, 266, 5, 273, 274] has been exceptional, exemplified by ChatGPT [303] developed by OpenAI. One key driver for the popularity and usability of recent LLMs is alignment. Alignment is a technique that aims to ensure that generated responses comply with human values. An example is illustrated in Figure 7. Currently, the standard procedure for aligning large language models (LLMs) primarily includes two approaches: SFT (Supervised Fine-Tuning) [221] and RLHF (Reinforcement Learning from Human Feedback) [55, 17]. Since LLMs have been used in a wide range of applications (e.g., editing/writing assistance, personal consultation, question answering, and customer support), many corresponding concerns would arise if the LLMs are not properly aligned otherwise.
Figure 7: An example of unaligned and aligned LLM response. The unaligned LLM response is problematic because it encourages risky financial behavior without considering the potential downsides. It fails to account for the ethical and responsible aspects of financial advice, potentially leading users to make harmful decisions.
@@
Characteristics of Alignment. To understand the characteristics of LLMs, a diverse array of benchmarks have been introduced [165, 284, 282, 283]. In contrast to general-purpose evaluation, alignment-focused evaluation depends on the taxonomy of alignment, associated with corresponding scenarios, criteria, and datasets [286, 350, 349]. Obtaining appropriate criteria and datasets for evaluating alignments in LLMs is crucial, albeit a non-trivial task  [51, 283]. This essentially involves representing the preferences of humans [41]. However, manually collecting human judgment can be expensive, time-consuming, and labor-intensive [345]. To address this issue, researchers proposed to use strong LLMs as an automated proxy for evaluating other LLMs [348]. For example, AUTO-J [154] is trained to tackle challenges in evaluating LLM alignments regarding generality, flexibility, and interoperability. AUTOCALIBRATE presents a multi-stage, gradient-free approach [154], to automatically calibrate and align an LLM-based evaluator toward human preference free of human intervention.
@@
Recent Advancements in Alignment. In the endeavor to align LLMs with human values, a myriad of research initiatives [263, 244, 257, 302, 313, 129, 248] have been undertaken to achieve effective LLM alignments. The forefront of these approaches emphasizes the generative capabilities of large language models (LLMs) for self-regulation with minimal human supervision. SELF-ALIGN [263] proposes a topic-guided, principle-driven approach to autonomously generate responses that are helpful, ethical, and reliable, leveraging the mechanism of in-context learning. Similarly, KNOWNO [244] is a framework for evaluating and aligning the uncertainty in LLM-based planning. Utilizing the theory of conformal prediction, KNOWNO ensures statistical reliability in task completion, thereby minimizing human assistance in complex planning scenarios. Additionally, PRO [257] introduces a response probability ranking method, enhancing the Bradley-Terry comparison model to effectively direct the LLM to favor the most appropriate response. Complementarily, P3O [302] presents a trajectory-wise policy gradient algorithm, which uniquely focuses on comparative rewards instead of traditional reward optimization trained from comparison-based losses.
@@
Discussion. The burgeoning field of LLM alignment, pivotal for the symbiosis of AI and humanity, anticipates transformative discoveries. Emphasizing the importance of AI safety and the seamless integration of AI with human society, prioritizing the alignment of LLMs, with human ethos is essential. As LLMs’ capabilities escalate, the complexity of achieving this alignment intensifies, necessitating increased scientific and technological investment. This demands an exploration of novel strategies in this domain. Foremost, amidst the rapid evolution of LLMs, it is crucial to guarantee their adherence to human ethical standards, which requires more theoretical breakthroughs [301]. In addition, the growing intricacy of AI architectures calls for automated systems capable of assessing and realigning these models [227]. Next, the black-box nature of LLMs also highlights the urgency for clarity and explainability in their alignment processes [343]. Lastly, leveraging adversarial attacks as a method to test and refine the alignment of LLMs emerges as an effective approach for ensuring their conformity to human values [355].
@@ 3.3 Law and Regulatory Compliance
Discussion. Despite the heroic striving of the AI Act to keep up with the accelerating dynamics of AI development, several discussions are also proposed around its practical compliance with LLMs. Hacker et al. argue that this direct regulation is unsatisfactory and could be further enhanced from 1) the definition of GPAIS, 2) the risk management of GPAIS, and 3) the adverse consequences for competition [95]. They propose to focus on the deployers and users more and directly apply non-discrimination and data protection law (GDPR compliance) on LLMs. Bommasani et al. [34] systematically evaluate the compliance with the draft EU AI Act of the foundation model providers like OpenAI and Google. They evaluate the compliance of 10 major foundation model providers (and their flagship models) with the 12 requirements proposed by the EU AI Act and use a scale from 0 (worst) to 4 (best) to rate each provider and model for each requirement. The best possible score for a provider or a model is 48, which indicates full compliance with the AI Act. Their results identify four areas where many organizations receive low scores (usually 0 or 1 out of 4) in terms of compliance with the AI Act: 1) copyrighted data, 2 compute/energy, 3) risk mitigation, and 4) evaluation/testing. Aside from these general regulations, there are also discussions on challenges of how to regulate LLMs for vertical domains such as medical usage [204] and healthcare [203].